# Learning in Strategic Queuing Systems with Small Buffers

### v1

@@ Abstract
Routers in networking use simple learning algorithms to find the best way to deliver packets to their desired destination. This simple, myopic and distributed decision system makes large queuing systems simple to operate, but at the same time, the system needs more capacity than would be required if all traffic were centrally coordinated. In a recent paper, Gaitonde and Tardos (EC 2020 and JACM 2023) initiate the study of such systems, modeling them as an infinitely repeated game in which routers compete for servers and the system maintains a state (number of packets held by each queue) that results from outcomes of previous rounds. Queues get to send a packet at each step to one of the servers, and servers attempt to process only one of the arriving packets, modeling routers. However, their model assumes that servers have no buffers at all, so queues have to resend all packets that were not served successfully. They show that, even with hugely increased server capacity relative to what is needed in the centrally-coordinated case, ensuring that the system is stable requires the use of timestamps and priority for older packets.
We consider a system with two important changes, which make the model more realistic: first we add a very small buffer to each server, allowing the server to hold on to a single packet to be served later (even if it fails to serve it); and second, we do not require timestamps or priority for older packets. Our main result is to show that when queues are learning, a small constant factor increase in server capacity, compared to what would be needed if centrally coordinating, suffices to keep the system stable, even if servers select randomly among packets arriving simultaneously.
This work contributes to the growing literature on the impact of selfish learning in systems with carryover effects between rounds: when outcomes in the present round affect the game in the future.
@@ 1 Introduction
In this paper, we model routers in networking as agents using simple learning algorithms to find the best way to get their packets served. This simple, myopic and distributed multi-agent decision system makes large queuing systems simple to operate, but at the same time, the system needs more capacity than would be required if all traffic were centrally coordinated. In a recent paper, Gaitonde and Tardos (2020, 2023), the authors initiated the study of such systems, modeling them as an infinitely repeated game in which routers compete for servers and the system maintains a state (number of packets held by each queue) that results from outcomes of previous rounds. Routers get to send a packet at each step at which they received a packet or have one in their queue [^1] The players are routers with states, the length of their queues; the terms router and queue are used interchangeably. to one of the servers, and each server attempts to process only one of the packets arriving to it. However, the model of Gaitonde and Tardos (2020, 2023) assumes that servers have no buffers at all, so queues have to resend all packets that were not served successfully. They show that, in their system, even with hugely increased server capacity relative to what is needed in the centrally-coordinated case, the use of timestamps and priority for older packets is required to ensure that the system is stable, i.e., that queue lengths do not diverge with time (see Section 2 for a formal definition).
@@
In networking systems, servers accepting packets typically have a very small buffer and can hold on to a few packets to be served later. In this paper, we consider the analog of the model of Gaitonde and Tardos (2020, 2023) in such systems with a tiny buffer at each server. We show that even with a buffer capable of holding only a single packet, the service capacity of the system greatly increases even when agents are learning. To see this, consider a system with a single queue receiving packets at a high rate, say with probability $\lambda=\frac{1}{2}$ at each time step, and multiple servers, each with a service rate (the probability of successfully serving a packet), say $\mu=\frac{1}{3}$. Without buffers, the condition for stability of such a system is that there is a single server with a service rate of $\mu>\lambda$, as a queue can only send a packet to one server at a time. So even with a large number of servers, each with service rate $\mu=\frac{1}{3}$, and thus, an arbitrarily large sum of service capacities, it would not be feasible to keep the system stable. Also, the learning algorithm of this single queue is facing a classical multi-arm bandit problem, aiming to learn which of the servers is the best one to send their packets to. By contrast, we observe that with a buffer of even just one packet at each server, a large enough number of servers with low capacity will always ensure stability, as the learning agent can now take advantage of low capacity servers as well, sending packets to them rarely enough such that they are likely to clear in time before the next packet is sent to them.
Back to our example of servers with $\mu=\frac{1}{3}$ and $\lambda=\frac{1}{2}$, which, as we saw above, is not stable without buffers for any number of servers, with buffers, two such servers will make the system stable if the queue alternates sending its packets to them. In light of this observation, the goal of a learning algorithm even with just one queue and many servers is now more complex: it should learn to distribute its packets in a way that takes advantage of the capacity of all the servers.
@@
Our main result, formally stated in Theorem 1, is to show that when multiple queues compete for service while using learning algorithms to identify good servers, a small constant-factor increase in server capacity—relative to what would be required under centralized coordination—suffices to maintain the system stable, even if servers randomly select among simultaneously arriving packets. Specifically, suppose we have $n$ queues with packet arrival rates of $\lambda_{1},\ldots,\lambda_{n}$, and $m$ servers of service rates $\mu_{1},\ldots,\mu_{m}$. Obviously, we must have $\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$, else the system cannot be stable even if fully coordinated. We show that this condition is not strong enough to ensure stability even with full coordination due to the small buffers. [^2] This condition allows to create a fractional assignment such that $\sum_{j}x_{ij}>\lambda_{i}$ and $\sum_{i}x_{ij}<\mu_{j}$. Using the matching decomposition of this assignment in a coordinated schedule will guarantee a bounded expected number of packets both at the queues and servers. However, our result shows that if $\lambda_{i}<\frac{1}{2}$ for all $i$, all queues use a learning algorithm that guarantees the no-regret condition with high probability, and $3\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$, this guarantees that the system remains stable. While we do not know whether the factor of 3 in the required capacity constraint is tight, we show that to guarantee that a no-regret outcome of learning by the queues remains stable, it is required to have at least $2\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$ (see Remark 1 and Theorem 3 for more details).
@@
The classical focus of work on scheduling in queuing systems is aimed at finding schedules that achieve optimal throughput (see, e.g., the textbook of Shortle et al. (2018)). For work evaluating efficiency loss due to selfishness in different classical queuing systems, see the book of Hassin Hassin (2020) and survey of Hassin and Haviv (2003). Closer to our motivation, there is a growing literature that aims to understand how systems perform when the queues use simple learning algorithm to find good service. Krishnasamy et al. (2016) considers a queue using a no-regret learning algorithm to find what may be the best servers, but does not consider multiple queues competing for service. Their primary goal is to study the expected size of the queue of packets as a function of time. They also extend the result to the case of multiple queues scheduled by a single coordinated scheduling algorithm, assuming there is a perfect matching between queues and optimal servers that can serve them. Sentenac et al. (2021) and Freund et al. (2022) extended this work to decentralized learning dynamics in bipartite queuing systems that attain near-optimal performance without the matching assumption. For a survey on the role of learning and information in queuing systems see Walton and Xu (2021).
@@
In our work, we focus on a game-theoretic model: queues using learning algorithms to best distribute their packets to get good service, while also competing for servers. We assume that each queue separately learns to selfishly make sure its own packets are served at the highest possible rate, offering a strategic model of scheduling packets in a queuing system. Closest to our model from this literature is the work Gaitonde and Tardos (2020, 2023), who consider the same bipartite model of queues and servers as we do, but without buffers. They show the exact condition to make such a system stable with central coordination of packets and prove that no-regret learning by the queues guarantees stability of the system if it has double the capacity needed for central coordination, assuming packets carry a timestamp and servers choose the oldest packet to serve. Fu et al. (2022) extended this work to a general network. Baudin et al. (2023) propose an alternative, episodic queuing system where agents have incentives to hold jobs in an episode before sending to a central server, but suffer penalties should their jobs not be completed before the end of the episode. They show that both equilibrium and no-regret outcomes ensure stability, so long as these costs are sufficiently large.
@@
In the works discussed so far, the servers receiving the packets have no buffers: all unserved packets are held in a queue and get resent to be served later. In queuing systems aimed at modeling networking, the receivers (servers) each do have a very small buffer, and can hold on to a couple of packets to be served later. It turns out that even having a buffer for a single packet significantly changes the effective service capacity of the network, as already explained above.
@@
In the case without such buffers, it was feasible to exactly characterize the capacity needed to be able to make the system stable with central coordination by an elegant use of linear programming duality (see Fu et al. (2022)). In the presence of limited buffer capacity at the servers, the rate at which a server accepts a packet depends on the state of the buffers. This is in contrast to a system without buffers, where the probability that a server accepts a packet is the service rate of the server, and is stable over time. There is some recent work considering queuing systems with such evolving service probabilities, see for example Grosof et al. (2024b, a). The goal of that work is to characterize the exact condition that allows the system to remain stable with central coordination of the schedule. Our work is the first to consider stability in such a system assuming each queue independently aims to optimize its own service rate running a no-regret learning algorithm.
@@
There is a large literature on bounding the price of anarchy for various games, beginning with Koutsoupias and Papadimitriou (1999). These results extend to analyzing outcomes in repeated games assuming players are no-regret learners Balcan et al. (2013); Roughgarden (2015); Syrgkanis and Tardos (2013). However, these works make the strong assumption that games in different rounds are independent. By contrast, different rounds in queuing games are not independent: the number of packets in the system depends on the success in previous periods. Work on games with carryover effects, or context, are considered in multi-agent reinforcement learning as well as in Markov games, see for example Littman (1994); Busoniu et al. (2008); Zhang et al. (2021). However, this line of work does not focus on the overall quality of learning outcomes. Another line of work on repeated games with such carryover effect between rounds, focusing on such overall learning outcomes, is the repeated ad-auction game with limited budgets, where the remaining budgets make the games no longer independent. See for example Balseiro and Gur (2019); Gaitonde et al. (2023); Fikioris and Tardos (2023); Fikioris et al. (2024).
@@ 2 Model
We consider a modified version of the model used by Gaitonde and Tardos (2020, 2023). There is a system of $n$ queues and $m$ servers. We divide time into discrete time steps $t=0,1,\dots$. In every step $t$, the following occurs. See the illustration in Figure 1.
@@
    Any packets not accepted into a server’s buffer are returned to their respective queues. Each queue receives bandit feedback based on the outcome of sending its packet: a reward of zero if the packet was not accepted by the server a reward of one if the packet was successfully placed in the server’s buffer. [^3] Feedback is received immediately after a packet is placed in the buffer, not necessarily when it is successfully processed.
Figure 1: A three-queue, three-server system in a single time step. Discs represent packets; left squares are queues, right squares are servers, and small squares are buffers. In this time step, two packets are sent to the top server, one of them is accepted into a buffer and then processed, the other one is returned to its queue. For the other two servers, one admits a packet into a buffer but fails to serve it and one processes a packet that was in its buffer.
@@
It is clear that if $\sum_{i}\lambda_{i}>\sum_{j}\mu_{j}$ then even a fully coordinated system cannot be stable, and the total number of packets waiting at the queues must grow linearly in time. In the case of a single queue with a single server, $\lambda<\mu$ is the necessary and sufficient condition for keeping the packets remaining in the system bounded by a constant (which depends on the gap in the inequality) at all times. The analogous condition is clearly necessary but not sufficient in case of multiple servers.
@@
Consider an example with 1 queue and 2 servers, each with $\mu=\frac{1}{2}$. We claim that $\lambda<\frac{23}{24}<1$ is needed for the system to be stable even with full coordination.
@@
To see why this is true, consider 3 consecutive time steps. To analyze the maximum processing rate we look at the scenario where the queue sends a packet in each of these steps. We claim that with probability $1/8$ it holds that in one of the three times the queue must fail to have a packet accepted by a server. To see this, assume the following.
  • Assume that at step 1 the queue sends a packet to one of the servers, which fails to serve it, but may keep it in its buffer. This event has probability at least $\frac{1}{2}$.
  • With full coordination, the queue is aware of the first buffer being full and can try to send the next packet to the second server. Assume that at this step neither of the two servers succeed in sending packets, but as before, may accept the new packet into its buffer. This event has probability at least $\frac{1}{4}$.
  • Even if the first two packets were accepted into buffers, in this case, at the third time step, both buffers are full, and so there is no place for the queue to send.
This proves that during a set of consecutive three steps the maximum expected service rate is at most
$\frac{\frac{1}{8}\cdot 2+\frac{7}{8}\cdot 3}{3}=\frac{23}{24}.$
Thus, if the arrival rate $\lambda$ is higher, we will have a buildup of the queue linear in $T$. ∎
@@ 3 Analysis
Our main result is the following theorem:
@@
Assuming $\sum_{i}\lambda_{i}<\frac{1}{3}\sum_{j}\mu_{j}$, $\lambda_{i}<\frac{1}{2}$ for all $i$, and all queues use a form of learning guaranteeing no-regret with high probability to identify servers they can use, the system remains stable with the expected number of packets in the system bounded by a (time-independent) constant at all times.
The key tool in our analysis is the following theorem of Pemantle and Rosenthal (1999):
@@
To use this theorem we will consider a long enough time in period of length $T$ and use a potential function that depends on the number of packets in each of the queues at the start of periods of length $T$. Assume that queue $i$ has $N_{i}$ packets at the start of the period. We will use the potential function $\Phi=\sum_{i}(N_{i}-(\lambda^{2}_{i}+3\delta)T)^{+}$, where $x^{+}=\max(x,0)$, and will aim to show that between the beginning and end of a $T$ long period the change in this potential satisfies the conditions of Theorem 1.
For a long enough period $T$, we expect that the packet arrivals and the servers’ service rates are all close enough to their expectations, and each queue has small enough regret in their learning algorithm. We will call a period of length $T$ goodδ for a parameter $\delta>0$ (to be chosen later) if the following three conditions are all satisfied. Later, we will also account for what happens when goodδ does not occur.
@@
The number of packets arriving to each queue $i$ in each period of length $\hat{T}\leq T$ is at most $\lambda_{i}\hat{T}+\delta T$.
@@
The learning algorithm of each queue $i$ accumulates regret at most $\delta T$ during the $T$-length period.
We use Chernoff and union bounds to bound from below the probability that conditions 1 and 2 hold when choosing a high enough $T$.
@@
To guarantee that Condition 3 holds with high probability, we will assume that queues use a learning algorithm with a high probability regret guarantee, such as EXP3.P.1, of Auer et al. (2002a), which offers a regret guarantee of $\sqrt{Tm\ln(mT/\eta)}$ with probability $1-\eta$.
@@
For any constants $\delta>0$ and $\eta>0$, by choosing the period length $T$ high enough we can guarantee that if all queues use EXP3.P.1 as their learning algorithm, the probability that Condition 3 holds for a single queue is at least $1-\eta$. Hence, the condition holds with probability $(1-n\eta).$
Now we are ready to start proving the main result. To show the expected decrease in the potential function, we consider whether a period of length $T$ satisfied the condition of being goodδ. By Lemmas 2 and 3, we see that the probability that goodδ fails to hold is bounded by $(m+2n)\eta$. The maximum possible increase in the potential function during a period of $T$ steps is at most $nT$ (if packets arrive during each step at each of the queues, and none reach the servers), so bad events contribute at most $(m+2n)\eta nT$ to the expected change in potential over the $T$ steps.
Next, we consider the change in potential for a goodδ period of length $T$.
Lemma 4.
If all servers have a packet in their buffer (possibly one that just arrived) for more than half of the iterations, and we are in the goodδ case, the total number of packets at the queues decreases by at least $(\frac{1}{2}\sum_{j}\mu_{j}-\sum_{i}\lambda_{i}-(n+m)\delta)T$.

### v2

@@ Abstract
We consider learning outcomes in games with carryover effects between rounds: when outcomes in the present round affect the game in the future. An important example of such systems is routers in networking, as they use simple learning algorithms to find the best way to deliver packets to their desired destination. This simple, myopic, and distributed decision process makes large queuing systems easy to operate, but at the same time, the system needs more capacity than would be required if all traffic were centrally coordinated. Gaitonde and Tardos (EC 2020 and JACM 2023) initiated the study of such systems, modeling them as an infinitely repeated game in which routers compete for servers and the system maintains a state (the number of packets held at each queue) that results from outcomes of previous rounds. However, their model assumes that servers have no buffers at all, so routers have to resend all packets that were not served successfully, which makes their system model unrealistic. They show that in their model, even with hugely increased server capacity relative to what is needed in the centrally coordinated case, ensuring that the system is stable requires the use of timestamps and priority for older packets.
We consider a system with two important changes, which make the model more realistic and allow for much higher traffic rates: first, we add a very small buffer to each server, allowing the server to hold on to a single packet to be served later (if it fails to serve it immediately), and second, we do not require timestamps or priority to older packets. Using theoretical analysis and simulations, we show that when queues are learning, a small constant-factor increase in server capacity, compared to what would be needed if centrally coordinating, suffices to keep the system stable, even if servers select randomly among packets arriving simultaneously.
@@ 1 Introduction
In this paper, we model routers in networking as agents using simple learning algorithms to find the best way to get their packets served. This simple, myopic, and distributed multi-agent decision system is an important application of multi-agent learning. Managing a networking system in such a distributed way makes large queuing systems simple to operate, but at the same time, the system needs more capacity than would be required if all traffic were centrally coordinated. In a recent paper, Gaitonde and Tardos (2020, 2023) initiated the study of such systems, modeling them as an infinitely repeated game in which routers compete for servers, and the system maintains a state (number of packets held by each queue) that results from outcomes of previous rounds. Routers get to send a packet at each step in which they received a packet or have one in their queue, [^1] Players are routers with states (their queue length); the terms router and queue are used interchangeably. to one of the servers. Each server attempts to process only one of the packets that arrive to it. However, the model of Gaitonde and Tardos (2020, 2023) (see also Sentenac et al. (2021); Freund et al. (2023)) assumes that servers have no buffers at all, so queues have to resend all packets that were not served successfully. They show that, in their system, even with hugely increased server capacity relative to what is needed in the centrally-coordinated case, the use of credible timestamps and priority for older packets by all servers is required to ensure that the system is stable, i.e., that queue lengths do not diverge with time (see Section 2 for a formal definition of stability).
@@
In networking systems, servers accepting packets typically have a very small buffer and can hold on to a few packets to be served later. In this paper, we consider the analog of the model of Gaitonde and Tardos (2020, 2023) in such systems with a tiny buffer at each server. We show that even with a buffer capable of holding only a single packet and non-coordinated learning agents, the service capacity of the system greatly increases. However, the tasks of the learning algorithms become more complex when servers have buffers. To see this, consider the case of a single queue with many possible servers. Without buffers, the queue faces a classical multi-arm bandit problem, aiming to learn which of the servers is the best one to send their packets to. By contrast, we observe that with a buffer of even just one packet at each server, a sufficiently large number of low-capacity servers will always ensure stability, as the learning agent can now take advantage of these low-capacity servers as well, sending packets to them rarely enough that they are likely to clear in time before the next packet is sent to them. This makes the goal of a learning algorithm more complex even with just one queue and many servers: it should learn to distribute its packets in a way that takes advantage of the capacity of all the servers. Interestingly, our results show that even simple bandit learning algorithms manage to learn to use low-capacity servers effectively in the presence of buffers.
@@
Our main result, formally stated in Theorem 1, is to show that when multiple agents compete for service while using learning algorithms to identify good servers, a small constant-factor increase in server capacity—relative to what would be required under centralized coordination—suffices to maintain the system stable, even if servers randomly select among simultaneously arriving packets. Specifically, suppose that we have $n$ queues with packet arrival rates of $\lambda_{1},\ldots,\lambda_{n}$, and $m$ servers with service rates $\mu_{1},\ldots,\mu_{m}$. Obviously, we must have $\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$, else the system cannot be stable even if fully coordinated. We show that this condition is not strong enough to ensure stability even with full coordination due to the small size of the buffers. [^2] We note that the condition allows to create a stable fully coordinated schedule, assuming we use large buffers at the servers. To see how to do this, note that the condition allows to create a fractional assignment such that $\sum_{j}x_{ij}>\lambda_{i}$ and $\sum_{i}x_{ij}<\mu_{j}$. Using the matching decomposition of this assignment in a coordinated schedule will guarantee a bounded expected number of packets both at the queues and at the servers. However, our result shows that if $\lambda_{i}<\frac{1}{2}$ for all $i$, all queues use any type of learning that achieves the no-regret condition with high probability, and $3\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$, then this guarantees that the system remains stable. While we do not know whether the factor of 3 in the required capacity constraint is tight and this remains as an open question, we provide a lower bound, showing that to guarantee that a no-regret outcome of learning by the queues remains stable, it is required to have at least $2\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$.
@@
The classical focus of work on scheduling in queuing systems is aimed at finding schedules that achieve optimal throughput (see, e.g., the textbook of Shortle et al. (2018)). For work evaluating efficiency loss due to selfishness in different classical queuing systems, see the book of Hassin (2020) and survey of Hassin and Haviv (2003). Closer to our motivation, there is a growing literature that aims to understand how systems perform when the queues use simple learning algorithm to find good service. Krishnasamy et al. (2016) considers a queue using a no-regret learning algorithm to find what may be the best servers, but does not consider multiple queues competing for service. Their primary goal is to study the expected size of the queue of packets as a function of time. They also extend the result to the case of multiple queues scheduled by a single coordinated scheduling algorithm, assuming there is a perfect matching between queues and optimal servers that can serve them. Sentenac et al. (2021) and Freund et al. (2022) extended this work to decentralized learning dynamics in bipartite queuing systems that attain near-optimal performance without the matching assumption. For a survey on the role of learning and information in queuing systems see Walton and Xu (2021).
@@
In our work, we diverge from the literature on queuing and scheduling and instead focus on a game-theoretic model with learning agents: queues using learning algorithms to best distribute their packets to get good service, while also repeatedly interacting with other learners and competing for servers. We assume that each queue separately learns to selfishly make sure its own packets are served at the highest possible rate, offering a strategic model of scheduling packets in a queuing system. Closest to our model from this literature is the work Gaitonde and Tardos (2020, 2023), who consider the same bipartite model of queues and servers as we do, but without buffers. They show the exact condition to make such a system stable with central coordination of packets and prove that no-regret learning by the queues guarantees stability of the system if it has double the capacity needed for central coordination, assuming packets carry a timestamp and servers choose the oldest packet to serve. Fu et al. (2022) extended this work to a general network. Baudin et al. (2023) propose an alternative, episodic queuing system where agents have incentives to hold jobs in an episode before sending to a central server, but suffer penalties should their jobs not be completed before the end of the episode. They show that both equilibrium and no-regret outcomes ensure stability, so long as these costs are sufficiently large.
@@
In the works discussed so far, the servers receiving the packets have no buffers: all unserved packets are held in a queue and get resent to be served later. In queuing systems aimed at modeling networking, the receivers (servers) each do have a very small buffer, and can hold on to a few packets to be served later. It turns out that even having a buffer for a single packet significantly changes the effective service capacity of the network with learning agents, as already explained above.
@@
In the case without such buffers, it was feasible to exactly characterize the capacity needed to be able to make the system stable with central coordination by an elegant use of linear programming duality (see Fu et al. (2022)). In the presence of limited buffer capacity at the servers, the rate at which a server accepts a packet depends on the state of the buffers. This is in contrast to a system without buffers, where the probability that a server accepts a packet is equal to the service rate of the server, and is stable over time. There is some recent work considering queuing systems with such evolving service probabilities, see for example Grosof et al. (2024b, a). The goal of that work is to characterize the exact condition that allows the system to remain stable with central coordination of the schedule. Our work is the first to consider stability in such a system assuming each queue independently aims to optimize its own service rate running a no-regret learning algorithm.
@@
There is a large literature on bounding the price of anarchy for various games, beginning with Koutsoupias and Papadimitriou (1999). These results extend to analyzing outcomes in repeated games assuming players are no-regret learners Balcan et al. (2013); Roughgarden (2015); Syrgkanis and Tardos (2013). However, these works make the strong assumption that games in different rounds are independent. By contrast, different rounds in queuing games are not independent: the number of packets in the system depends on the success in previous periods. Work on games with carryover effects, or context, are considered in multi-agent reinforcement learning as well as in Markov games, see for example Littman (1994); Busoniu et al. (2008); Zhang et al. (2021). However, this line of work does not focus on the overall quality of learning outcomes. Another line of work on repeated games with such carryover effects between rounds, focusing on such overall learning outcomes, is the repeated ad-auction game with limited budgets, where the remaining budgets make the games no longer independent. See for example Balseiro and Gur (2019); Gaitonde et al. (2023); Fikioris and Tardos (2023); Fikioris et al. (2024).
@@ 2 Model
We consider a more realistic version of the model used by Gaitonde and Tardos (2020, 2023). There is a system of $n$ queues and $m$ servers. We divide time into discrete time steps $t=0,1,\dots$, where in every step $t$, the following occurs. See also the illustration in Figure 1.
@@
    Any packets not accepted into a server’s buffer are returned to their respective queues. Each queue receives bandit feedback based on the outcome of sending its packet: a reward of 00 if the packet was not accepted by the server or $1$ if the packet was successfully placed in the server’s buffer. [^3] Feedback is immediate after a packet is placed in the buffer, not necessarily when it is successfully processed.
Figure 1: A system with three queues and three servers in a single time step. Discs represent packets; the left squares are queues, the right squares are servers, and the small squares are buffers. In this time step, two packets are sent to the top server. One of these packets is accepted into the buffer and then processed, while the other is returned to its queue. For the other two servers, one admits a packet into its buffer but fails to serve it, and the other processes a packet that was already in its buffer from a previous round.
@@
It is clear that if $\sum_{i}\lambda_{i}>\sum_{j}\mu_{j}$ then even a fully coordinated system cannot be stable, and the total number of packets waiting at the queues must grow linearly in time. In the case of a single queue with a single server, $\lambda<\mu$ is the necessary and sufficient condition for keeping the packets remaining in the system bounded by a constant (which depends on the gap in the inequality) at all times; see, e.g., Shortle et al. (2018). The analogous condition in the case of multiple servers is clearly necessary but not sufficient.
@@
Consider an example with one queue and two servers, each with service probability $\mu=\frac{1}{2}$. We claim that $\lambda<\frac{23}{24}<1$ is needed for the system to be stable even with full coordination.
@@
To see why this is true, consider 3 consecutive time steps. To analyze the maximum processing rate, we look at the scenario where the queue sends a packet in each of these steps. We claim that with probability at least $1/8$ it holds that in one of the three times, the queue must fail to have a packet accepted by a server. Assume that the packet sent in the first of the three periods is accepted into a buffer but not processed for two steps. This event has probability at least $\frac{1}{4}$. With coordination, a packet in the next period will be sent to the other server. Assume that this packet is not processed when it is sent, which has probability at least $\frac{1}{2}$. Now in the third period, both buffers are full, so no packet can be accepted.
This proves that during three consecutive steps, the maximum expected service rate is at most $\frac{\frac{1}{8}\cdot 2+\frac{7}{8}\cdot 3}{3}=\frac{23}{24}.$ Thus, if the arrival rate $\lambda$ is higher, we have a buildup of the queue linear in $T$. ∎
@@ 3 Analysis
Our main result is the following theorem. This result holds under either bandit or full feedback, and applies to arbitrary learning algorithms, as long as they achieve sublinear regret.
@@
Assuming $\sum_{i}\lambda_{i}<\frac{1}{3}\sum_{j}\mu_{j}$, $\lambda_{i}<\frac{1}{2}$ for all $i$, and all queues use a form of learning guaranteeing low regret with high probability to identify servers they can use, the system remains stable with the expected number of packets in the system bounded by a (time-independent) constant at all times.
The key tool in our analysis is the following theorem of Pemantle and Rosenthal (1999). Informally, the theorem states that a stochastic sequence that has negative drift when it grows large and is sufficiently regular (has bounded moments) is bounded by a constant at all times.
@@
To use this theorem, we consider a long enough time period of length $T$ and use a potential function that depends on the number of packets in each of the queues at the start of that time period. Assume that queue $i$ has $N_{i}$ packets at the start of the period. We will use the potential function $\Phi=\sum_{i}(N_{i}-(\frac{1}{2}\lambda_{i}+2\delta)T)^{+}$, where $x^{+}=\max(x,0)$, and will aim to show that between the beginning and end of a $T$ long period, the change in this potential satisfies the conditions of Theorem 1.
For $T$ large enough, we expect the empirical packet arrivals and servers’ service rates are all close enough to their expectations, and each queue has small enough regret for their learning algorithm. We will call a period of length $T$ goodδ for a parameter $\delta>0$ (to be chosen later) if the following three conditions are all satisfied. Later, we will also account for what happens in the “bad event” when goodδ does not occur.
@@
The number of packets arriving at each queue $i$ during each period of length $\hat{T}\leq T$ is at most $\lambda_{i}\hat{T}+\delta T$.
@@
The learning algorithm of each queue $i$ accumulates regret at most $\delta T$ during the $T$-length period for the decisions it made during this period when it had a packet to send.
To show the expected decrease in the potential function, we consider whether a period of length $T$ satisfied the condition of being goodδ. We will show below (in Lemmas 4 and 5) that the probability that goodδ fails to hold is bounded by $(m+2n)\eta$ for a small $\eta$, as shown by the two lemmas. The maximum possible increase in the potential function during a period of $T$ steps is at most $nT$ (if packets arrive during each step at each of the queues, and none reach the servers), so bad events contribute at most $(m+2n)\eta nT$ to the expected change in potential over the $T$ steps.
To show the decrease in the potential in a goodδ period of length $T$, the following two lemmas separately consider servers that serve packets most of the time and those that do not.
Lemma 2.
If all servers have a packet in their buffer (possibly one that just arrived) for more than half of the iterations, and we are in the goodδ case, then the total number of packets at the queues decreases by at least $(\frac{1}{2}\sum_{j}\mu_{j}-\sum_{i}\lambda_{i}-(n+m)\delta)T$.
Proof.
By Condition 2, the total number of packets served is at least $\sum_{j}(\frac{1}{2}\mu_{j}-\delta)T$. By Condition 1, the total number of arriving packets is at most $\sum_{i}(\lambda_{i}+\delta)T$. Combining these two bounds establishes the lemma. ∎
Lemma 3.
Suppose there is a server whose buffer is empty more than half of the $T$ iterations, and we are in the $\textit{good}_{\delta}$ case. Consider a queue $i$ and assume that $\delta<\frac{1}{2}(\frac{1}{2}-\lambda_{i})$. Then, either the number of packets at this queue decreases during the period of length $T$, or the number of packets left in it at the end of the period is at most $N^{\prime}_{i}\leq(\frac{1}{2}\lambda_{i}+2\delta)T$. If at the start of the period there are at least $T$ packets in the queue, then the number decreases by at least $(\frac{1}{2}-\lambda_{i}-2\delta)T$.
Proof.
If the queue has packets in every step in the window, by the no-regret Condition 3, it clears at least $(\frac{1}{2}-\delta)T$. By Condition 1, it receives at most $(\lambda_{i}+\delta)T$ packets, so the number of packets goes down by at least $(\frac{1}{2}-\lambda_{i}-2\delta)T$. This happens, in particular, when the queue has at least $T$ packets at the beginning of the window.
Now suppose the queue has packets in the last $\hat{T}$ steps, but was empty in the previous step. If $\hat{T}\leq\frac{1}{2}T$, then by Condition 1, in the last $\hat{T}$ steps it receives at most $(\frac{1}{2}\lambda_{i}+\delta)T$ new packets. So $N^{\prime}_{i}\leq(\frac{1}{2}\lambda_{i}+\delta)T$.
The rest of the proof focuses on the case when $\hat{T}>\frac{1}{2}T$. Let $I$ be the number of steps queue $i$ and the server were empty. So the server was empty at least $\frac{1}{2}T-I$ steps that queue $i$ was attempting to send packets. By the no-regret Condition 3, queue $i$ must have succeeded in clearing close to this many packets. Let $S$ be the number of packets queue $i$ received before the last $\hat{T}$ steps. Before the last $\hat{T}$ steps, queue $i$ had and cleared $N_{i}+S$ packets, so it was empty at most $T-\hat{T}-N_{i}-S$ steps, and this is also an upper bound for $I$. Using this upper bound with the no-regret Condition 3, we get that queue $i$ cleared at least $\frac{1}{2}T-I-\delta T\geq\hat{T}-(\frac{1}{2}+\delta)T+N_{i}+S$ packets. Queue $i$ had in total $N_{i}+S$ packets before the last $\hat{T}$ steps, where by Condition 1 it received at most $\lambda_{i}\hat{T}+\delta T$ packets. Therefore:
$N_{i}^{\prime}$$\leq$ $\left(N_{i}+S+\lambda_{i}\hat{T}+\delta T\right)-\left(\hat{T}-\left(\frac{1}{2}+\delta\right)T+N_{i}+S\right)$
$=$ $\left(\frac{1}{2}+2\delta\right)T+(\lambda_{i}-1)\hat{T}\leq\left(\frac{1}{2}+2\delta\right)T+\frac{1}{2}(\lambda_{i}-1)T=\left(\frac{1}{2}\lambda_{i}+2\delta\right)T,$
where the last inequality follows from $\lambda_{i}<1$ and $\hat{T}>\frac{1}{2}T$. ∎
To prove Theorem 1 we now need to bound the probability that goodδ fails to hold. We use Chernoff and union bounds to bound from below the probability that conditions 1 and 2 hold when choosing a sufficiently large $T$.
@@
To guarantee that Condition 3 holds with high probability, we will assume that queues use learning algorithms that have high-probability regret guarantees: Let $R_{T}$ be the regret of a queue. We say that a queue has low regret with high probability if for every $\eta>0$, there exists a sublinear bound $\epsilon(T)=o(T)$ such that $Pr[R_{T}\leq\epsilon(T)]>1-\eta$. This can be guaranteed, for example, using the algorithm EXP3.P.1 of Auer et al. (2002a).
@@
Suppose that all the queues use learning methods that guarantee low regret with high probability. Then, for every $\delta>0$ and $\eta>0$, by choosing the period length $T$ high enough we can guarantee that the probability that Condition 3 holds for a single queue is at least $1-\eta$. Hence, the condition holds with probability at least $(1-n\eta).$
We are now ready to start proving the main result.